# Cross-Modal Consistency Learning for Sign Language Recognition

### v1

@@ Abstract
Pre-training has been proven to be effective in boosting the performance of Isolated Sign Language Recognition (ISLR). Existing pre-training methods solely focus on the compact pose data, which eliminate background perturbation but inevitably suffer from insufficient semantic cues compared to raw RGB videos. Nevertheless, direct representation learning only from RGB videos remains challenging due to the presence of sign-independent visual features. To address this dilemma, we propose a Cross-modal Consistency Learning framework (CCL-SLR), which leverages the cross-modal consistency from both RGB and pose modalities based on self-supervised pre-training. First, CCL-SLR employs contrastive learning for instance discrimination within and across modalities. Through the single-modal and cross-modal contrastive learning, CCL-SLR gradually aligns the feature spaces of RGB and pose modalities, thereby extracting consistent sign representations. Second, we further introduce Motion-Preserving Masking (MPM) and Semantic Positive Mining (SPM) techniques to improve cross-modal consistency from the perspective of data augmentation and sample similarity, respectively. Extensive experiments on four ISLR benchmarks show that CCL-SLR achieves impressive performance, demonstrating its effectiveness. The code will be released to the public.
@@ 1 Introduction
Sign language, which combines body gestures and facial expressions to convey semantic meaning, serves as the primary communication tool within the Deaf community. To facilitate a barrier-free understanding between sign language and spoken language, extensive research [17, 44, 28, 21, 55, 13, 16, 46, 48] has explored sign language understanding from various perspectives. Isolated Sign Language Recognition (ISLR) is a fundamental task in sign language understanding, which focuses on recognizing the meaning of sign language videos at the word level to achieve fine-grained understanding. For the deployment of ISLR models in real-world applications, developing robust visual encoders with enhanced representation capabilities has increasingly been recognized as a crucial research direction.
@@
Current ISLR methods mainly employ supervised paradigms on small-scale dataset [26, 24, 18, 15], which can be categorized into two paradigms based on the input modalities. (1) RGB-based methods [17, 44, 28, 26, 27, 25, 24, 14, 15, 29] explore various strategies to extract discriminative visual features from RGB videos. Although these methods are intuitive and straightforward, they fail to explicitly capture key gesture-related information. Furthermore, signer-dependent variations hinders effective feature extraction, ultimately compromising the robustness of the sign representations. (2) Pose-assisted methods [21, 20, 55] integrate the pose modality and employ multi-stream fusion frameworks to leverage cross-modal complementary information. By utilizing the compact pose modality, these methods efficiently eliminate non-skeletal noise. However, current pose data is directly derived from pose extractors trained on general domain data distributions. Due to motion blur and hand occlusion in sign language videos, these general extractors inevitably produce errors and missing detection, limiting the performance ceiling of these pose-assisted methods.
@@
Although these supervised methods have achieved promising results, the scarcity of labeled training data inherently constrains their representation and generalization capabilities. To tackle this problem, recent works [13, 16, 46, 48, 47] attempt to explore self-supervised pre-training strategies in the pose modality to enhance visual modeling capabilities. For instance, MASA [47] propose a motion-aware masked auto encoder to utilizes the pose modality by introducing a motion prior and global semantic cues into pre-training. However, it only incorporate RGB information by late fusion in fine-tuning, failing to leverage original RGB features during pre-training. As shown in Fig. 1(a), these self-supervised methods focus solely on extracting visual information from pose modality. This paradigm fails to consider cross-modal feature interactions and overlooks the potential benefits of original RGB visual information.
@@
In this paper, we aim to introduce cross-modal consistency learning into the pre-training paradigm to fully enhance feature extraction in subsequent stages. Notably, effectively applying self-supervised paradigms such as contrastive learning to cross-modal interaction is non-trivial due to two key challenges: (1) Redundant information in the RGB modality. RGB images contain redundant non-semantic information in background regions, such as signer-dependent variations such as body type and skin color, which are often irrelevant to pose information. These redundant cues interfere with the feature extraction of sign language-related cues, further impacting the ISLR performance. (2) Inconsistency between modalities. Although the RGB and pose data of the same instance correspond to each other, the similarity structural relationships between instances across the two modalities are inconsistent, which leads to the inherent misalignment in feature distributions across RGB and pose modalities.
@@
To this end, we propose a novel Cross-modal Consistency Learning framework (CCL-SLR), which aims to model consistency constraints during pre-training, as shown in Fig. 1 (b). We employ single-modal contrastive learning and cross-modal contrastive learning to facilitate instance discrimination within individual modalities and across different modalities, respectively. To more effectively extract consistent sign language representations from both RGB and pose modalities in the contrastive learning paradigm, our approach incorporates two key modules: Motion-Preserving Masking (MPM) and Semantic Positive Mining (SPM). (1) To eliminate impact from redundant features, MPM introduces an off-the-shelf flow-based model to extract latent representations, leveraging this information to capture motion dynamics. By preserving motion-related information, MPM suppresses non-semantic regions in RGB videos, thereby enhancing the relevance of learned features. (2) To capture cross-modal structural consistency, SPM identifies semantically similar samples within each modality to provide pseudo-supervision for both modalities. Through a multi-label classification approach, this pseudo-supervision introduces shared positive constraints across different modalities, effectively bridging the distribution gap between them. To validate the effectiveness of our method, we conduct extensive experiments on four widely adopted ISLR benchmarks. In summary, our contributions are as follows:
@@
  • We propose a novel pre-training framework named CCL-SLR, which effectively leverages the single-modal and cross-modal contrasrive learning on both RGB and pose modality for enhancing ISLR.
@@
  • We introduce MPM and SPM, two innovative techniques designed to mitigate the distribution gap between modalities and further strengthen the consistency constrains during cross-modal pre-traininng.
@@
  • We conduct extensive experiments on four widely-adopted benchmarks and achieve impressive performance, which demonstrates effectiveness and representation capbility of CCL-SLR.
@@ 2 Related Work
Figure 2: (a) The CCL-SLR pre-training framework. RGB and pose sequences are processed through two single-modal branches to learn modality-specific representation. During this process, MPM is applied to mask motion-irrelevant regions in RGB videos. Additionally, cross-modal positive samples are identified through SPM, which are utilized for $\mathcal{L}^{R}$ and $\mathcal{L}^{L}$. Finally, the consistency of RGB and pose embedding spaces is constrained through $\mathcal{L}_{cross}$. (b) An overview of SPM. SPM selects the top-k most similar samples from each modality-specific memory bank and merges them to form pseudo-labels.
@@
Sign Language Recognition. Sign language recognition is a subtask in field of sign language understanding, which aims to recognize single sign word within a given video. Early works [6, 7, 37, 32, 34] utilize hand-craft feature to model spatio-temporal representation. With development of deep convolution neural networks (CNNs), numerous recent works [26, 27, 25, 24, 14, 15, 29] leverage CNN-based architectures such as TSM [30], R3D [10] and I3D [1] to extract visual representation from RGB videos. For example, StepNet [36] use TSM to model part-level spatial and temporal relationship, achieving impressive performance on ISLR.
@@
However, these RGB-based methods may suffer from static irrelevant cues in sign language videos [55], such as background and clothing. Consequently, several works begin to explore pose-assisted method, trying to utilize the effectiveness of compact sign pose data. For instance, NLA-SLR [55] present a video-keypoint multi-stream backbone and utilize semantic information within sign glosses to promote ISLR. SAM-v2 [21] jointly leverage pre-extracted multi-modal data with a branching ensemble model to achieve higher recognition rate. Although these works effectively fuse RGB and pose modality features, they rely on complex model designs and feature fusion strategies, which affect their scalability. Meanwhile, some studies [13, 16, 47, 46, 48] incorporate self-supervised pre-training paradigms into ISLR to address the challenge of data scarcity. SignBERT [13] enhance model representation capacity through generative pretext task specifically designed for pose data during self-supervised pre-training. BEST [46] leverage discrete variational autoencoder to compress hand shapes into a codebook for implicit clustering, serving as pseudo-labels to better learn skeletal priors. Nevertheless, these approaches overlook the benefits of aligning consistency visual representations of RGB and pose data during the pre-training for downstream ISLR tasks.
@@
In this work, we focus on ISLR pre-training paradigm, learning instance discriminative representations through contrastive learning within and across modalities.
@@
Contrastive Representation Learning. Contrastive learning has significantly advanced self-supervised visual representation learning by performing instance discrimination in a fully self-supervised manner [11, 3, 2, 43, 8, 33]. It pulls positive samples closer and pushes negative samples away. Since there are no supervised labels, different data augmentations of the same instance are treated as a positive pair, while samples from other instances are used as negative pairs. Recently, MoCo [11] and MoCo v2 [3] strengthens contrastive learning by maintaining a dynamic memory bank of negative sample embeddings from previous iterations. Based on them, current advances have demonstrated substantial improvements in representation capability through various technical innovations, including positive mining [54, 52, 42, 19] and data augmentation [45, 5, 38, 53].
@@
Although these methods have achieved success in various downstream tasks, incorporating the contrastive learning paradigm into multi-modal ISLR pre-training remains a challenge. To this end, we propose CCL-SLR, which attempts to extend the MoCo v2 framework to the challenging task of cross-modal sign language representation learning.
@@ 3 Method
As illustrated in Fig. 2, our framework first applies the modality-specific data augmentation to generate augmented samples. Subsequently, these augmented samples are processed through their respective encoders for feature extraction. The encoder outputs are then utilized to compute the contrastive loss, facilitating the learning of instance discriminative representations within and across modalities. For the RGB modality, we introduce a novel Motion-Preserving Masking mechanism to mitigate redundant non-semantic visual cues in the RGB video. Furthermore, to enhance the single-modal discriminative representation learning while leveraging the cross-modal consistency, we propose a Semantic Positive Mining mechanism that identifies semantically similar positive sample pairs.
@@ 3.1 Preliminaries
Single-modal contrastive learning has emerged as a powerful paradigm in diverse tasks, with notable success demonstrated by MoCo v2 [3]. For a given single-modal input $x$ in a mini-batch, the learning process begins with data augmentation to generate two distinct samples $x_{q}$ and $x_{k}$ (query and key). Then, two encoders, $\mathcal{E}_{q}$ and $\mathcal{E}_{k}$, are applied to embed $x_{q}$ and $x_{k}$ into a hidden space:
@@
where $\mathcal{E}_{q}$ and $\mathcal{E}_{k}$ denote the query and key encoders, $\theta_{q}$ and $\theta_{k}$ denote the learnable parameters of the encoders. Notably, the key encoder $\mathcal{E}_{k}$ is maintained as an exponential moving average (EMA) of the query encoder $\mathcal{E}_{q}$, updated as: $\theta_{k}\leftarrow\gamma\theta_{k}+(1-\gamma)\theta_{q}$, where $\gamma$ denotes the momentum coefficient. During the pre-training phase, instance discrimination is achieved through the InfoNCE [39] loss:
@@ 3.2 Single-Modal Contrastive Learning
Framework Architecture. As illustrated in Fig. 2, CCL-SLR consists of two parallel single-modal branches: an RGB branch and a pose branch. Each branch encompasses two pathways for key and query sample, along with two queue-based memory banks. Each pathway consists of an encoder and a projection head. Since recent ISLR datasets do not provide pose annotation, we first extract the corresponding pose sequence $X^{P}\in\mathbb{R}^{T\times J\times C_{p}}$ using an off-the-shelf estimator [4] from RGB frame sequence $X^{R}\in\mathbb{R}^{T\times H\times W\times 3}$. Here, $T$ denotes the frame length, $J$ denotes the number of joints, while $H$ and $W$ correspond to the spatial dimensions. $C_{p}$ represents XY coordinates with confidence of pose data.
First, we use spatial-temporal augmentation to generate different query and key sample for single-modal branches, denoted as $X^{m}_{s}$, where $m\in\{R,P\}$ represents RGB and pose modality and $s\in\{q,k\}$ represents query and key samples. Second, each sample $X^{m}_{s}$ is processed by its corresponding encoder $\mathcal{E}^{m}_{s}$ to obtain embedding ${g}^{m}_{s}\in\mathbb{R}^{d}$. The RGB encoder employs R3D-50 [10] backbone, while the pose encoder utilizes a GCN+Transformer architecture [50]. Then, ${g}^{m}_{s}$ is projected using multi-layer perceptron (MLP) projection head $\mathcal{P}^{m}_{s}$ to obtain the final embedding $f^{m}_{s}\in\mathbb{R}^{d}$. Two queue-based memory banks $\mathcal{M}^{m}_{g}\in\mathbb{R}^{N\times d}$ and $\mathcal{M}^{m}_{f}\in\mathbb{R}^{N\times d}$ are utilized for storing L2 normalized embeddings of ${g}^{m}_{k}$ and ${f}^{m}_{k}$, respectively:
@@
Figure 3: An overview of MPM. Here $h(\cdot)$ and $h^{-1}(\cdot)$ denotes the transformation from pixel/latent space to latent/pixel space, respectively. The original video is encoded into latent space using a flow-based generative model. Channels in the latent representation with low standard deviation along temporal dimension are then masked out. The masked latent representation is decoded back into pixel space. The reconstructed frames are then processed through gray-scaling and binarization to produce motion-aware mask sequence.
Motion-Preserving Masking (MPM). Unlike pose sequence that focuses on keypoint coordinates, RGB video contains substantial non-gestural information such as background and clothing. To preserve essential motion components while suppressing irrelevant regions, we propose MPM, an augmentation that enhances motion capture capabilities and mitigates bias from non-skeleton elements. For mapping RGB videos from pixel space to a compact feature space, MPM employs a pre-trained flow-based generative model, AdvFlow [31] for motion-preserving latent representation extraction.
Given an RGB video $X^{R}$, we encode it into latent representation $Z\in\mathbb{R}^{T\times d_{z}}$. Empirically, the temporal standard deviation of each channel serves as an indicator of its motion information content. We apply this measure to $Z$ to preserve dynamic features while suppressing static ones:
@@
where $\mathds{1}(\cdot)$ denotes the indicator function, $\sigma(\cdot)$ computes the standard deviation along the temporal channel, and $p$ controls the mask intensity. The motion-preserving video $\hat{X}^{R}\in\mathbb{R}^{T\times H\times W\times 3}$ is subsequently reconstructed from the masked latent representation $\hat{Z}$. We then derive the binary mask sequence $\zeta\in\mathbb{R}^{T\times H\times W\times 1}$ from $\hat{X}^{R}$ and apply it to $X^{R}$ to generate the augmented sample.
Semantic Positive Mining (SPM). While single-modal contrastive learning enables CCL-SLR to extract modality-specific representation, the “semantic collision” problem identified by CrosSCLR [54] remains a significant challenge. The semantic collision occurs during contrastive learning when sample embeddings with high semantic similarity in the memory bank are mistakenly treated as negative examples. In response to this semantic misalignment, we introduce SPM, a mechanism that identifies semantic positive samples across memory banks of different modalities. As shown in Fig. 2 (b), we select the top-$k$ nearest neighbors of ${g}^{m}_{q}$ from $\mathcal{M}^{m}_{g}$ as positive samples, while treating the remaining embeddings as negative samples. This selection process generates a binary pseudo-label $y^{m}\in\mathbb{R}^{N}$:
@@
To align the neighborhood structure across modalities, we combine $y^{R}$ and $y^{P}$ through maximization to obtain the cross-modal binary pseudo-label $y$. Actually, since the labels are binary, the max function here is equivalent to taking the union of the positive samples from two modalities. The embeddings in $\mathcal{M}^{m}_{f}$ serve as classifiers, generating class logits $\hat{y}^{m}\in\mathbb{R}^{N}$ that classify ${f}^{m}_{q}$ into $N$ pseudo categories:
@@
where $\tau_{spm}$ denotes the temperature coefficient and $\phi(\cdot)$ is the sigmoid function [51]. Then, $\hat{y}^{m}$ and $y$ are utilized for binary cross-entropy (BCE) loss:
@@
The final single-modal contrastive loss is as follows:

### v2

@@ Abstract
Pre-training has been proven to be effective in boosting the performance of Isolated Sign Language Recognition (ISLR). Existing pre-training methods solely focus on the compact pose data, which eliminates background perturbation but inevitably suffers from insufficient semantic cues compared to raw RGB videos. Nevertheless, learning representation directly from RGB videos remains challenging due to the presence of sign-irrelevant visual features. To address this dilemma, we propose a Cross-modal Consistency Learning framework (CCL-SLR), which leverages cross-modal consistency between both RGB and pose modalities in a self-supervised paradigm. First, CCL-SLR employs contrastive learning for instance discrimination within and across modalities. Through single-modal and cross-modal contrastive learning, CCL-SLR gradually aligns the feature spaces of RGB and pose modalities, thereby extracting consistency sign representation. Second, we further introduce Motion-Preserving Masking (MPM) and Semantic Positive Mining (SPM) techniques to improve cross-modal consistency from the perspective of data augmentation and sample clustering, respectively. Extensive experiments on four ISLR benchmarks show that CCL-SLR achieves impressive performance, demonstrating its effectiveness. The code will be released to the public.
@@ 1 Introduction
Sign language, which combines body gestures and facial expressions to convey semantic meaning, serves as the primary communication tool within the Deaf community. To facilitate a barrier-free understanding between sign language and spoken language, extensive research [17, 44, 28, 21, 55, 13, 16, 46, 48] has explored sign language understanding from various perspectives. Isolated Sign Language Recognition (ISLR) is a fundamental task in sign language understanding, that focuses on recognizing the meaning of sign language videos at the word level to achieve fine-grained understanding. For the deployment of ISLR models in real-world applications, developing robust visual encoders with enhanced representation capabilities has increasingly been recognized as a crucial research direction.
@@
Current ISLR methods mainly employ supervised paradigms on small-scale datasets [26, 24, 18, 15], which can be categorized into two paradigms based on the input modalities. (1) RGB-based methods [28, 26, 27, 25, 24, 14, 15, 29] explore various strategies to extract discriminative visual features from RGB videos. Although these methods are intuitive and straightforward, they fail to explicitly capture key gesture-related information. Furthermore, signer-dependent variations hinder effective feature extraction, ultimately compromising the robustness of the sign representations. (2) Pose-assisted methods [17, 44, 21, 20, 55] integrate the pose modality and employ multi-stream fusion frameworks to leverage cross-modal complementary information. By utilizing the compact pose modality, these methods efficiently eliminate non-skeletal noise. However, current pose data is directly derived from general-domain pose extractors. Due to motion blur and hand occlusion in sign language videos, these general-domain extractors inevitably produce errors and missing detections, limiting the performance ceiling of these pose-assisted methods.
@@
Although these supervised methods have achieved promising results, the scarcity of labeled training data inherently constrains their representation and generalization capabilities. To tackle this problem, recent works [13, 16, 46, 48, 47] attempt to explore self-supervised pre-training strategies in the pose modality to enhance visual modeling capabilities. For instance, MASA [47] proposes a motion-aware masked autoencoder to utilize the pose modality by introducing a motion prior and global semantic cues into pre-training. However, it incorporates only RGB information through late fusion during fine-tuning, failing to leverage original RGB features during pre-training. As shown in Fig. 1(a), these self-supervised methods focus solely on extracting visual information from pose modality. This paradigm fails to consider cross-modal feature interactions and overlooks the potential benefits of original RGB visual information.
@@
In this paper, we aim to introduce cross-modal consistency learning into the pre-training paradigm to fully enhance feature extraction in subsequent stages. Notably, effectively applying self-supervised paradigms such as contrastive learning to cross-modal interaction is non-trivial due to two key challenges: (1) Redundant information in the RGB modality. RGB images contain redundant non-semantic information in background regions, such as signer-dependent variations such as body type and skin color, which are often irrelevant to pose information. These redundant cues disrupt the extraction of sign language-related features, further impacting the ISLR performance. (2) Inconsistency between modalities. Although the RGB and pose data of the same instance correspond to each other, the similarity of structural relationships between instances across the two modalities are inconsistent, leading to the inherent misalignment in feature distributions across RGB and pose modalities.
@@
To this end, we propose a novel Cross-modal Consistency Learning framework (CCL-SLR), which aims to model consistency constraints during pre-training, as shown in Fig. 1 (b). We employ single-modal contrastive learning and cross-modal contrastive learning to facilitate instance discrimination within individual modalities and across different modalities, respectively. To more effectively extract consistent sign language representations from both RGB and pose modalities in the contrastive learning paradigm, our approach incorporates two key modules: Motion-Preserving Masking (MPM) and Semantic Positive Mining (SPM). (1) To eliminate impact from redundant features, MPM introduces an off-the-shelf flow-based model [31] to extract latent representations, leveraging this information to capture motion dynamics. By preserving motion-related information, MPM suppresses non-semantic regions in RGB videos, thereby enhancing the relevance of learned features. (2) To capture cross-modal structural consistency, SPM identifies semantically similar samples within each modality to provide pseudo-supervision for both modalities. Through a multi-label classification approach, this pseudo-supervision introduces shared positive constraints across different modalities, effectively bridging the distribution gap between them. To validate the effectiveness of our method, we conduct extensive experiments on four widely adopted ISLR benchmarks. In summary, our contributions are as follows:
@@
  • We propose a novel pre-training framework named CCL-SLR, which effectively leverages the single-modal and cross-modal contrastive learning on both RGB and pose modalities for enhancing ISLR.
@@
  • We introduce MPM and SPM, two innovative techniques designed to mitigate the distribution gap between modalities and further strengthen the consistency constraints during cross-modal pre-training.
@@
  • We conduct extensive experiments on four widely-adopted benchmarks and achieve impressive performance, which demonstrates effectiveness and representation capability of CCL-SLR.
@@ 2 Related Work
Figure 2: (a) The CCL-SLR pre-training framework. RGB and pose sequences are processed through two single-modal branches to learn modality-specific representation. During this process, MPM is applied to mask motion-irrelevant regions in the RGB videos. Additionally, cross-modal positive samples are identified through SPM, which are utilized for single-modal loss $\mathcal{L}^{R}$ and $\mathcal{L}^{L}$. Finally, the consistency of RGB and pose embedding spaces is constrained through cross-modal contrastive loss $\mathcal{L}_{cross}$. (b) An overview of the Semantic Positive Mining. SPM selects the top-$k$ most similar samples from each modality-specific memory bank and merges them to get pseudo-labels.
@@
Sign Language Recognition. Sign language recognition is a subtask in the field of sign language understanding, which aims to recognize a single sign word within a given video. Early works [6, 7, 37, 32, 34] utilize hand-crafted features to model spatio-temporal representation. With the development of deep convolution neural networks (CNNs), numerous recent works [26, 27, 25, 24, 14, 15, 29] leverage CNN-based architectures such as TSM [30], R3D [10], and I3D [1] to extract visual representations from RGB videos. For example, StepNet [36] uses TSM to model part-level spatial and temporal relationships, achieving impressive performance on ISLR.
@@
However, these RGB-based methods may suffer from static irrelevant cues in sign language videos [55], such as background and clothing. Consequently, several works have begun exploring pose-assisted methods, leveraging the compact and informative nature of sign pose data. For instance, NLA-SLR [55] presents a video-keypoint multi-stream backbone and utilize semantic information within sign glosses to promote ISLR. SAM-v2 [21] jointly leverages pre-extracted multi-modal data with a branching ensemble model to achieve a higher recognition rate. Although these works effectively fuse RGB and pose modality-specific features, they rely on complex model designs and feature fusion strategies, which limits their scalability. Meanwhile, some studies [13, 16, 47, 46, 48] incorporate self-supervised pre-training paradigms into ISLR to address the challenge of data scarcity. SignBERT [13] enhances model representation capacity through a generative pretext task specifically designed for pose data during self-supervised pre-training. BEST [46] leverage discrete variational autoencoder to compress hand shapes into a codebook for implicit clustering, serving as pseudo-labels to better learn skeletal priors. Nevertheless, these approaches overlook the benefits of aligning consistency visual representations of RGB and pose data during pre-training for downstream ISLR tasks.
@@
In this work, we focus on the ISLR pre-training paradigm, learning instance discriminative representations through contrastive learning within and across modalities.
@@
Contrastive Representation Learning. Contrastive learning has significantly advanced self-supervised visual representation learning by performing instance discrimination in a fully self-supervised manner [11, 3, 2, 43, 8, 33]. It pulls positive samples closer and pushes negative samples away. Since there are no supervised labels, different data augmentations of the same instance are treated as a positive pair, while samples from other instances are used as negative pairs. Recently, MoCo [11] and MoCo v2 [3] have strengthened contrastive learning by maintaining a dynamic memory bank of negative sample embedding from previous iterations. Building on these methods, current advances have demonstrated substantial improvements in representation capability through various technical innovations, including positive mining [54, 52, 42, 19] and data augmentation [45, 5, 38, 53].
@@
Despite their success in various downstream tasks, integrating the contrastive learning paradigm into multi-modal ISLR pre-training remains challenging. To this end, we propose CCL-SLR, which extends the MoCo v2 framework to the challenging task of multi-modal sign language representation learning.
@@ 3 Method
As illustrated in Fig. 2, our framework first applies modality-specific data augmentation to generate augmented samples. Subsequently, these augmented samples are fed into their respective encoders for feature extraction. The encoder outputs are then utilized to compute the contrastive loss, facilitating the learning of instance discriminative representations within and across modalities. For the RGB modality, we introduce a novel Motion-Preserving Masking technique to suppress redundant non-semantic visual cues in RGB videos. Furthermore, to enhance single-modal representation learning while leveraging cross-modal consistency, we propose a Semantic Positive Mining strategy that identifies semantically similar positive pairs.
@@ 3.1 Preliminaries
Since CCL-SLR involves two modalities, we first introduce contrastive learning in the single-modal case to simplify the explanation. Single-modal contrastive learning has emerged as a powerful paradigm across various tasks, with notable success demonstrated by MoCo v2 [3]. Given a single-modal input $x$ in a mini-batch, the learning process begins with data augmentation to generate two distinct samples $x_{q}$ and $x_{k}$ (query and key). Two encoders, $\mathcal{E}_{q}$ and $\mathcal{E}_{k}$, are then applied to embed $x_{q}$ and $x_{k}$ into a hidden space:
@@
where $\mathcal{E}_{q}$ and $\mathcal{E}_{k}$ denote the query and key encoders, $\theta_{q}$ and $\theta_{k}$ denote the learnable parameters of the encoders. Notably, the key encoder $\mathcal{E}_{k}$ is maintained as an exponential moving average (EMA) of the query encoder $\mathcal{E}_{q}$, updated as:
$\theta_{k}\leftarrow\gamma\theta_{k}+(1-\gamma)\theta_{q},$ (2)
where $\gamma$ denotes the momentum coefficient. During the pre-training phase, instance discrimination is achieved through the InfoNCE [39] loss:
@@ 3.2 Single-Modal Contrastive Learning
Framework Architecture. As illustrated in Fig. 2, CCL-SLR consists of two parallel single-modal branches: an RGB branch and a pose branch. Each branch encompasses two pathways for key and query samples, along with two queue-based memory banks. Each pathway consists of an encoder and a projection head. We extract the corresponding pose sequence $X^{P}\in\mathbb{R}^{T\times J\times C_{p}}$ from the RGB frame sequence $X^{R}\in\mathbb{R}^{T\times H\times W\times 3}$ using an off-the-shelf estimator [4]. Here, $T$ denotes the frame length, $J$ denotes the number of joints, and $H$ and $W$ correspond to the spatial dimensions. $C_{p}=3$ represents the XY coordinates along with the confidence score of pose data.
First, we use spatial-temporal augmentation to generate different query and key samples for each single-modal branch, denoted as $X^{m}_{s}$, where $m\in\{R,P\}$ represents RGB and pose modalities and $s\in\{q,k\}$ represents query and key samples. Second, each sample $X^{m}_{s}$ passes through its corresponding encoder $\mathcal{E}^{m}_{s}$ to obtain an embedding ${g}^{m}_{s}\in\mathbb{R}^{d}$. The RGB encoder employs R3D-50 [10] backbone, while the pose encoder utilizes a GCN+Transformer architecture [50]. Then, ${g}^{m}_{s}$ is projected using a multi-layer perceptron (MLP) projection head $\mathcal{P}^{m}_{s}$ to obtain the final embedding $f^{m}_{s}\in\mathbb{R}^{d}$. Two queue-based memory banks $\mathcal{M}^{m}_{g}\in\mathbb{R}^{N\times d}$ and $\mathcal{M}^{m}_{f}\in\mathbb{R}^{N\times d}$ are utilized for storing L2-normalized embeddings of ${g}^{m}_{k}$ and ${f}^{m}_{k}$, respectively:
@@
Figure 3: An overview of MPM. $h(\cdot)$ and $h^{-1}(\cdot)$ denote the transformation from pixel/latent space to latent/pixel space, respectively. The video is encoded into latent space using a flow-based generative model. Channels in the latent representation with low standard deviation along the temporal dimension are then masked out. The masked latent representation is decoded back into pixel space. The reconstructed frames are then processed through gray-scaling and binarization to produce motion-aware mask sequence.
Motion-Preserving Masking (MPM). Unlike pose sequences, which focus on keypoint coordinates, RGB videos contain substantial non-gestural information such as background and clothing. To preserve essential motion components while suppressing irrelevant regions, we propose MPM, a data augmentation technique that enhances motion capture capabilities and mitigates bias from non-skeletal elements. As shown in Fig.3, MPM employs a pre-trained flow-based generative model, AdvFlow [31], to transform original RGB videos into motion-preserving videos and generate a mask sequence for data augmentation.
Concretely, given an RGB video $X^{R}$, we encode it into a latent representation $Z\in\mathbb{R}^{T\times d_{z}}$. Empirically, the temporal standard deviation of each channel reliably indicates its motion content. We leverage this measure on $Z$ to preserve dynamic features while suppressing static ones:
@@
where $\mathds{1}(\cdot)$ denotes the indicator function, $\otimes$ denotes the element-wise multiplication, $\sigma(\cdot)$ computes the standard deviation along the temporal channel, and $p$ controls the mask intensity. The motion-preserving video $\hat{X}^{R}\in\mathbb{R}^{T\times H\times W\times 3}$ is then decoded from the masked latent representation $\hat{Z}$. Subsequently, we derive the binary mask sequence $\zeta\in\mathbb{R}^{T\times H\times W}$ from $\hat{X}^{R}$ and apply it to $X^{R}$ with probability $\alpha$ to generate the augmented sample.
Semantic Positive Mining (SPM). While single-modal contrastive learning enables CCL-SLR to extract modality-specific representations, the “semantic collision” problem identified by CrosSCLR [54] remains a challenge. The semantic collision occurs during contrastive learning when sample embeddings with high semantic similarity in the memory bank are mistakenly treated as negative examples. To mitigate this semantic misalignment, we introduce SPM, a mechanism that identifies semantic positive samples across memory banks of different modalities. As illustrated in Fig. 2 (b), we select the top-$k$ nearest neighbors of ${g}^{m}_{q}$ from $\mathcal{M}^{m}_{g}$ as positive samples, while treating the remaining embeddings as negative samples. This selection process generates binary pseudo-labels $y^{m}\in\mathbb{R}^{N}$ for both modalities:
@@
To align the neighborhood structure across modalities, we combine $y^{R}$ and $y^{P}$ by taking their element-wise maximum to obtain the cross-modal binary pseudo-label $y$. The embeddings in $\mathcal{M}^{m}_{f}$ serve as classifiers, generating logits $\hat{y}^{m}\in\mathbb{R}^{N}$ that classify ${f}^{m}_{q}$ into $N$ pseudo categories:
@@
where $\tau_{spm}$ denotes the temperature coefficient and $\phi(\cdot)$ is the sigmoid function [51]. Then, $\hat{y}^{m}$ and $y$ are utilized for binary cross-entropy (BCE) loss to help single-modal contrastive learning bring the query sample closer to semantic positive samples from both modalities:
@@
The final single-modal loss is formulated as::